Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipt was $824 and twice as many adult tickets as student tickets were sold then how many of each were sold?
Answer by mananth(16946)

Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipt was $824 and twice as many adult tickets as student tickets were sold then how many of each were sold?
let the number of students be x
The number of adults be 2x
2x+3*2x= 824
2x+6x=824
8x=824
x=824/8
x=103 which is the number of students
The number of adults will be 2*103= 206